OpenStudy (anonymous):

Line H is represented by the following equation: 2x + 2y = 8 What is most likely the equation for line K so the set of equations has infinitely many solutions? x + 2y = 2 x + y = 4 2x + y = 6 x − y = 8

OpenStudy (texaschic101):

infinitely many solutions....you are gonna need to find a line that is the same for there to be infinite solutions. That means it has to have the same slope and the same y intercept. So you need to put everything in y = mx + b form and compare the slopes and the y intercepts.

OpenStudy (texaschic101):

2x + 2y = 8 2y = -2x + 8 y = -x + 4 In y = mx + b form, the slope is in the m position and the y intercept is in the b position. So the slope is -1 and the y intercept is 4. Now put your answer choices in y = mx + b form....your answer will have a slope of -1 and a y intercept of 4. Any ideas now ?

OpenStudy (texaschic101):

or an easier way would be to reduce 2x + 2y = 8 by dividing every term by 2.....leaving you with what ?

OpenStudy (anonymous):

so its the last one im pretty sure because you said to divide everyone by one. So it leaves me with the last one?

OpenStudy (texaschic101):

no... 2x/2 + 2y/2 = 8/2 x + y = 4 understand ? This equation, along with the one given, will result in infinite solutions

OpenStudy (anonymous):

Haha i didnt want the answer i wanted to try and solve it myself... im a little slow in math :P but thank you for trying to help me :)

OpenStudy (texaschic101):

do you understand what I did ? Do you have any questions at all ?

OpenStudy (anonymous):

Yes i kinda did im still a bit confused but i get the concept thank you :)